If A and B are independent events, then P .A and B/ D P.A/  P .B/


1
Expert's answer
2021-07-16T08:37:06-0400

We know that,

If A and B are independent events then,

"P(A/B)=P(A)\\P(B/A)=P(B)\ \ and\\P(A \ and \ B)=P(A)\cdot P(B)"


So, from the third equation we can say that,


"\dfrac{P(A\ and\ B)}{P(A)\times P(B)}=1"